### v1

@@ IV Combating Beam Squint in Non-TTD-based Hybrid Beamformers
The high cost and implementation complexity of TTD units have motivated the development of alternative beam squint mitigation strategies for hybrid beamformers that avoid the use of TTDs. For instance, the authors of [4] recently proposed a Non-TTD-based wideband hybrid beamforming scheme based on a new metric, the wideband beam gain (WBBG), which leverages conventional phase shifters. However, the performance of this approach has not been compared or analyzed against TTD-based solutions. This section outlines the core concept of the Non-TTD-WBBG method and discusses its strengths and limitations relative to TTD-based architectures.
Consider a transmitter equipped with $N$ antenna elements transmitting multi-carrier signals over a wide frequency band spanning from $f^{\mathrm{min}}$ to $f^{\mathrm{max}}$, as illustrated in Fig. 2-a, where the intended user is located at the angular position $\phi_{\mathrm{u}}$ and is expected to receive signals from all carriers. When conventional narrowband maximum-ratio transmission (MRT) beamforming is employed using frequency-independent (wideband) phase shifters, the signals at different carrier frequencies are steered toward different directions, deviating from $\phi_{\mathrm{u}}$. This frequency-dependent deviation results in an angular spread ranging from $\phi_{\mathrm{u}}^{\mathrm{min}}$ to $\phi_{\mathrm{u}}^{\mathrm{max}}$, corresponding to $f^{\mathrm{min}}$ and $f^{\mathrm{max}}$, respectively, as depicted in Fig. 2-b. This beam squint effect causes each carrier’s beam to be radiated towards a distinct angle, often placing the user outside the main lobe of many carriers. A Full-TTD architecture can counteract beam squint by aligning all frequency components toward $\phi_{\mathrm{u}}$, thereby achieving the full array gain of $N$ for every carrier, as illustrated in the figure. In contrast, the Non-TTD-WBBG approach proposed in [4] mitigates beam squint using conventional phase shifters. This method optimizes the phase shifter values to maximize the minimum beam gain across all carriers within the angular interval $[\phi_{\mathrm{u}}^{\mathrm{min}},\phi_{\mathrm{u}}^{\mathrm{max}}]$, resulting in a relatively flat wideband beam pattern, as shown in Fig. 2-c. Unlike conventional MRT, the WBBG technique provides approximately uniform beam gain across all carriers, thus ensuring max-min fairness under beam squint.
@@
  • Optimizing the max-min fairness criterion for large-scale antenna arrays across a wide range of carrier frequencies introduces substantial computational complexity, particularly in the hybrid mode for jointly designing the digital precoders and analog beamformers. This complexity is generally much higher than addressing hybrid beamforming involving only max or min objectives. In practice, the problem is often addressed through a combination of convex approximations and iterative algorithms such as bisection methods [4]. However, these approximations can lead to a notable performance gap compared to optimal baselines that do not enforce max-min fairness constraints.
@@
  • Similar to the Sparse-TTD architecture discussed earlier, the Non-TTD-WBBG scheme requires wideband phase shifters, the effective deployment of which poses significant challenges in terms of cost and efficiency in wideband and ultra-wideband communication scenarios.
@@
In summary, Non-TTD-WBBG provides a viable alternative for moderate bandwidths but experiences performance degradation in wideband and ultra-wideband scenarios. Moreover, its cost-effectiveness is uncertain due to the complexities of implementing wideband phase shifters, as discussed in the previous section.
@@
V Optimal Wideband Hybrid Beamformers: A Cost-Performance Trade-off Analysis
@@
Figure 3: Performance versus fractional bandwidth for various beamforming architectures. We consider $N=128$ antennas, $M=16$, $\tilde{\gamma}=0$ dB, and $\theta_{\mathrm{u}}=60^{\circ}$. A simple linear model is used for the $\mathcal{E}(b)$, with unity gain at $b=0$ and a maximum degradation of 6 dB at $b=\pm\frac{B}{2f_{0}}$. Depending on the design parameters and fractional bandwidth, the Sparse-TTD-NBBG beamformer’s achieved performance lies at some point within the upper and lower bounds shown as the green-shaded region.
@@
Figure 4: Cost trend versus desired operating fractional bandwidth for various beamforming architectures. The regions where Non-TTD-WBBG, Sparse-TTD-NBBG, and Full-TTD-NBBG beamformers are more cost-effective are highlighted in orange, blue, and green, respectively.
@@
As discussed earlier, the fractional bandwidth $B/f_{0}$ plays a pivotal role in shaping both the performance and cost of wideband hybrid beamformers, making the selection of an optimal architecture a non-trivial design challenge. In this section, we present a cost-performance analysis using simplified system models to provide insights into this trade-off. Fig. 3 compares the performance of several hybrid beamforming architectures: Non-TTD wideband beam-gain (Non-TTD-WBBG), Non-TTD narrowband beam-gain (Non-TTD-NBBG), Full-TTD-NBBG, and Sparse-TTD-NBBG, all evaluated as functions of the fractional bandwidth $B/f_{0}$. Unlike the Non-TTD-WBBG beamformer, which explicitly broadens the beam across the entire spectrum, the other architectures optimize performance without explicitly considering inter-carrier frequency interactions or beam widening. Therefore, to distinguish these approaches against WBBG, we have used the suffix NBBG representing a narrowband beam-gain approach for these structures.
To simplify the analysis, we ignore the bandwidth-dependent performance degradation of TTDs by assuming ideal units that provide frequency-independent delays, as they are typically designed for wideband operation. While the Full-TTD-NBBG architecture can, in theory, eliminate beam squint entirely, making its performance nearly independent of the fractional bandwidth, the practical limitations of wideband phase shifters in Sparse-TTD and Non-TTD architectures introduce performance degradation as the operating bandwidth increases. To quantify this, we consider a simplified performance model. Assume an ideal line-of-sight channel with a transmitter employing an $N$-element half-wavelength ULA, transmitting over $12M+$ 1 uniformly spaced carrier frequencies across bandwidth $B$ with central frequency $f_{0}$. Let the fractional baseband carrier vector be defined as $\boldsymbol{b}=\left[-\frac{B}{2f_{0}},\ldots,\frac{B}{2f_{0}}\right]^{\mathrm{T}}_{(2M+1)\times 1}$, and let $G_{\mathrm{NB}}(b,\theta_{\mathrm{u}})$ denote the array gain at fractional baseband carrier $b$ and observation angle $\theta_{\mathrm{u}}$.
First, we examine the Non-TTD narrowband beamforming approach. The sum-SE performance gap between the Full-TTD-NBBG and Non-TTD-NBBG beamformers denoted by $\Delta R_{\mathrm{FN-NN}}=R_{\mathrm{FN}}-R_{\mathrm{NN}}$ can be approximated by
@@
$\Delta R_{\mathrm{FN-NN}}=\sum_{m=-M}^{M}\log\left(\frac{1+\tilde{\gamma}N\eta}{1+\tilde{\gamma}G(b_{m},\theta_{\mathrm{u}})\mathcal{E}(b_{m})\eta}\right)$ (2)
@@
where $\tilde{\gamma}$ is the SNR at each antenna element, [^1] Under an ideal far-field line-of-sight model, the SNR is identical across all antenna elements. $\mathcal{E}(b)\leq 1$ is a decreasing function of $b$ that models the impact of frequency-dependent hardware impairments in the phase shifters evaluated at $f=f_{0}(1+b)$, and $\eta<1$ captures the performance degradation due to hybrid beamforming relative to a fully digital implementation. The numerator of the logarithmic term includes the maximum array gain $N$ optimally realized by the Full-TTD structure, while the denominator includes the gain $G(b_{m},\theta_{\mathrm{u}})=G_{\mathrm{NB}}(b_{m},\theta_{\mathrm{u}})$ achieved by the NBBG beamformer for each carrier $b_{m}$. This gain can be approximated by the array gain of a fully digital beamformer with ideal phase shifters, given by $G_{\mathrm{NB}}(b,\theta_{\mathrm{u}})\approx\left|\frac{\sin(N\pi\Delta_{b})}{\sin(\pi\Delta_{b})}\right|$, where $\Delta_{b}=\frac{b}{2}\sin(\theta_{\mathrm{u}})$. This leads to the performance gap $\Delta R_{\mathrm{FN-NN}}$ shown in Fig. 3. It is observed that the gap becomes significantly large for $B/f_{0}>0.05$, rendering this architecture suitable only for narrowband scenarios. The Non-TTD-WBBG beamformer mitigates this gap by optimizing a wideband beam pattern $G_{\mathrm{WB}}(b_{m},\theta_{\mathrm{u}})$ that remains approximately flat across all carriers $b_{m}$ via a max-min beam gain optimization. This optimization is a complex non-convex problem which can be approximately solved using convexification techniques, such as those in [14]. Once the beam pattern is optimized through calculating the optimal value of $G_{\mathrm{WB}}(b_{m},\theta_{\mathrm{u}}),\forall m$, the corresponding performance gap between the Full-TTD-NBBG and Non-TTD-WBBG is obtained as $\Delta R_{\mathrm{FN-NW}}=R_{\mathrm{FN}}-R_{\mathrm{NW}}$ obtained from (2) where $G(b_{m},\theta_{\mathrm{u}})=G_{\mathrm{WB}}(b_{m},\theta_{\mathrm{u}})$. As depicted in Fig. 3, the Non-TTD-WBBG architecture substantially reduces the performance gap compared to the Non-TTD-NBBG approach. However, the gap increases with fractional bandwidth due to two main factors: the degradation of wideband beam gain caused by the increased angular coverage required to accommodate squinted beams, and the worsening hardware impairments of phase shifters at higher fractional bandwidths.
@@
Now we discuss the performance of the Sparse-TTD-NBBG beamformer. Depending on the structure of the hybrid beamformer and the number of RF chains and TTDs employed for the Sparse-TTD architecture, the performance can be analyzed to lie between two bounds. The lower bound corresponds to the case where no TTDs are used, which can be approximated by the performance of the Non-TTD-WBBG architecture. The upper bound is achieved when the maximum number of TTDs (i.e., $N$) is utilized together with $N$ available phase shifters in a fully digital beamformer structure, under the assumption that phase shifters are still present, unlike in the Full-TTD configuration. This corresponds to allocating each antenna element a dedicated TTD cascaded with a phase shifter. Since the cascade of a TTD and a phase shifter is functionally equivalent to a TTD, the performance of this configuration is similar to that of Full-TTD, but with additional degradation due to the bandwidth-dependent impairments of the phase shifters. Accordingly, the performance upper bound can be approximated as $\sum_{m=-M}^{M}\log\left(1+\tilde{\gamma}N\mathcal{E}(b_{m})\right)$. Depending on the design parameters, the actual performance of the Sparse-TTD-NBBG beamformer will lie somewhere between these bounds, as illustrated by the green-shaded region in Fig. 3.
@@
Beyond performance, practical deployment also necessitates consideration of implementation cost. Due to the wide range of parameters influencing cost, the literature lacks a definitive comparison of the cost across different wideband hybrid beamforming architectures. However, estimated comparisons can be made based on qualitative cost considerations and implementation assumptions. Assuming similar baseband and RF chains across all architectures represented in Fig. 1, the primary cost differences arise from the use of TTD units and wideband phase shifters. Among the studied architectures, the Non-TTD-NBBG offers the lowest cost but suffers from severe performance degradation at higher fractional bandwidths, limiting its suitability to narrowband systems. The cost trend with respect to the desired operational fractional bandwidth for the other beamforming architectures is schematically illustrated in Fig. 4. Sparse-TTD architecture commonly incurs higher costs than the Non-TTD counterparts due to the additional expense of TTD units. However, this cost gap narrows as the fractional bandwidth increases, since wideband phase shifters, employed in both Sparse-TTD and Non-TTD-WBBG architectures, become increasingly expensive at higher fractional bandwidths, with their cost growth rate eventually surpassing that of TTDs. More specifically, in ultra-wideband scenarios, the cost of Sparse-TTD and Non-TTD-WBBG architectures may ultimately exceed that of the Full-TTD architecture as illustrated in Fig. 4. This is due to the increased complexity of compensating components, such as wideband NGD circuits and calibration mechanisms, required to maintain constant phase for wideband phase shifters at extreme bandwidths, as discussed in Section III.
@@
Finally, an overall cost-performance analysis can be conducted to guide the selection of the optimal architecture as illustrated in Fig. 5. Based on the preceding discussion, the Non-TTD-NBBG and Full-TTD-NBBG hybrid beamformers clearly emerge as the preferred choices for narrowband and ultra-wideband scenarios, respectively. More specifically, the Full-TTD-NBBG beamformer is the preferred option in ultra-wideband scenarios in terms of both performance and cost as shown in Figs. 3 and 4, respectively, and highlighted in Fig. 5. In contrast, the preferred option for wideband scenarios can be Sparse-TTD-NBBG or Non-TTD-WBBG; however, the decision between the optimal architecture is more nuanced, as it depends on specific system design parameters and the relative emphasis placed on performance versus implementation cost.

### v2

@@ IV Combating Beam Squint in Non-TTD-based Hybrid Beamformers
The high cost and implementation complexity of TTD units have motivated the development of alternative beam squint mitigation strategies for hybrid beamformers that avoid the use of TTDs. For instance, the authors of [10] recently proposed a Non-TTD-based wideband hybrid beamforming scheme based on a new metric, the wideband beam gain (WBBG), which leverages conventional PSs. However, the performance of this approach has not been compared or analyzed against TTD-based solutions. This section outlines the core concept of the Non-TTD-WBBG method and discusses its strengths and limitations relative to TTD-based architectures.
Consider a transmitter equipped with $N$ antenna elements transmitting multi-carrier signals over a wide frequency band spanning from $f^{\mathrm{min}}$ to $f^{\mathrm{max}}$, as illustrated in Fig. 3-a, where the intended user is located at the angular position $\phi_{\mathrm{u}}$ and is expected to receive signals from all carriers. When conventional narrowband maximum-ratio transmission (MRT) beamforming is employed using frequency-independent (wideband) PSs, the signals at different carrier frequencies are steered toward different directions, deviating from $\phi_{\mathrm{u}}$. This frequency-dependent deviation results in an angular spread ranging from $\phi_{\mathrm{u}}^{\mathrm{min}}$ to $\phi_{\mathrm{u}}^{\mathrm{max}}$, corresponding to $f^{\mathrm{min}}$ and $f^{\mathrm{max}}$, respectively, as depicted in Fig. 3-b. This beam squint effect causes each carrier’s beam to be radiated towards a distinct angle, often placing the user outside the main lobe of many carriers. A Full-TTD architecture can counteract beam squint by aligning all frequency components toward $\phi_{\mathrm{u}}$, thereby achieving the full array gain of $N$ for every carrier, as illustrated in the figure. In contrast, the Non-TTD-WBBG approach proposed in [10] mitigates beam squint using conventional PSs. This method optimizes the PS values to maximize the minimum beam gain across all carriers within the angular interval $\phi_{\mathrm{u}}^{\mathrm{min}},\phi_{\mathrm{u}}^{\mathrm{max}}]$, resulting in a relatively flat wideband beam pattern, as shown in Fig. 3-c. Unlike conventional MRT, the WBBG technique provides approximately uniform beam gain across all carriers, thus ensuring max-min fairness under beam squint.
@@
  • Optimizing the max-min fairness criterion for large-scale antenna arrays across a wide range of carrier frequencies introduces substantial computational complexity, particularly in the hybrid mode for jointly designing the digital precoders and analog beamformers. This complexity is generally much higher than addressing hybrid beamforming involving only max or min objectives. In practice, the problem is often addressed through a combination of convex approximations and iterative algorithms such as bisection methods [10]. However, these approximations can lead to a notable performance gaps.
@@
  • Similar to the Sparse-TTD architecture discussed earlier, the Non-TTD-WBBG scheme requires wideband PSs, the effective deployment of which poses significant challenges in terms of cost and efficiency in wideband and ultra-wideband communication scenarios.
@@
In summary, Non-TTD-WBBG provides a viable alternative for moderate bandwidths but experiences performance degradation in wideband and ultra-wideband scenarios. Moreover, its cost-effectiveness is uncertain due to the complexities of implementing wideband PSs, as discussed before.
V Challenges and Considerations in Wideband TTD/PS Implementation
This section examines the practical implementation challenges of wideband PSs and TTDs across the three aforementioned beamforming architectures. Technical specifications and key limitations for several commercial and lab-fabricated implementations, ranging from lower frequency bands to the sub-THz regime, are summarized in Table I.
TABLE I: Representative Wideband PSs and TTDs and Key Associated Loss/Error Values
Device Resolution Freq. Range Time Delay Range Max Phase/Delay Error Insertion Loss
Commercial Digitally Controlled Wideband PSs
PS-100M20G-10B-SFF-MIL-SM (Quantic PMI) 10 Bit 0.1–20 GHz N/A $\leq\pm 6^{\circ}$ 0-13 dB
TGP2105-SM (Qorvo) 6-bit 6–18 GHz N/A $\leq\pm 8^{\circ}$ 6-10 dB
Commercial Digitally Controlled Wideband TTDs
ADAR4002 TDU (Analog Devices) 7 Bit 0.5–19 GHz 254 ps $\leq\ 2$ ps 1-20 dB
MM7005T (Miller MMIC) 4 Bit 1–12 GHz 390 ps $\leq\ 10$ ps 6-9 dB
Commercial Sub-THz Digitally Controlled Wideband PS and TTD
529 Series (Mi-Wave) 4 Bit 75–110 GHz N/A $\leq\pm 4^{\circ}$ 0-3 dB
https://doi.org/10.23919/EuMC48046.2021.9338128 continuous 53–120 GHz N/A N/A 0-3 dB
V-A Wideband PSs
From a hardware realization standpoint, wideband and ultra-wideband beamforming architectures are fundamentally constrained by implementation-level impairments in digitally controlled PSs. Unlike idealized frequency-flat models, commercial wideband PSs exhibit measurable frequency-dependent phase distortion and frequency-dependent insertion loss across their operating bandwidth. For example, as seen in Table I, the Qorvo TGP2105-SM operating over 6–18 GHz reports a frequency-dependent phase error of $8^{\circ},+8^{\circ}]$ and an insertion loss ranging from 6 to 10 dB across the whole bandwidth. These distortions stem from dispersion in transmission-line networks, reactive elements employed in switched filter cells, and parasitic effects introduced by semiconductor switching networks. For small fractional bandwidths, these impairments remain moderate, rendering PS-based architectures compact and cost-effective with negligible amplitude/phase distortion.
V-B Wideband TTDs
In contrast to the frequency-dependent phase distortion of wideband PSs, commercial wideband TTDs aim to enforce frequency-invariant delay and exhibit less severe frequency dependent insertion loss, but introduce a different set of hardware challenges; they typically exhibit delay-dependent insertion loss and nonzero delay error across the operating band due to switching nonidealities and dispersion in delay lines. As shown in Table I, representative devices such as the ADAR4002 TDU and MM7005T demonstrate delay errors on the order of a few picoseconds over wide bandwidths. Note that although commercial TTDs introduce delay-dependent insertion loss and signal distortions, compensating for these impairments is generally less challenging than mitigating the frequency-dependent distortions inherent to wideband PSs. Since TTDs required delays are initially computed by the beamformer and applied to each TTD, amplifiers with gains adjusted proportionally to the applied calculated delays can be readily employed to alleviate associated loss and distortion.
V-C Other Impairments and Sub-THz Considerations
Beyond standard frequency-dependent phase, delay, and amplitude distortions, wideband electrical PSs and TTDs suffer from electrically large parasitics, high dielectric losses, and extreme sensitivity to fabrication tolerances. These factors are further compounded by thermal fluctuations, which modify material properties and physical dimensions as well as device bias points, inducing time-varying phase errors and cumulative delay drift. While complex multi-section electrical architectures can attempt to equalize these responses, they incur excessive power consumption and calibration overhead. These implementation challenges become significantly more pronounced at sub-THz frequencies, where ohmic and dielectric losses scale aggressively, parasitic reactances become dominant, and structural fabrication margins represent a substantial fraction of the operating wavelength. Photonic architectures provide a robust alternative to mitigate these electrical limitations. At sub-THz frequencies, photonic wideband TTDs are significantly simpler to implement than PS counterparts because frequency-independent delays arise naturally from optical path lengths or dispersion. In contrast, photonic PSs require stringent wavelength, thermal, and fabrication tolerances to achieve stable wideband operation. The final two rows of Table I present sub-THz wideband commercial electrical wideband PS and photonic TTD, respectively.
@@ V-C Other Impairments and Sub-THz Considerations
VI Optimal Wideband Hybrid Beamformers: A Cost-Performance Trade-off Analysis
(a) Refer to caption (b)
@@ VI Optimal Wideband Hybrid Beamformers: A Cost-Performance Trade-off Analysis
Figure 4: (a): Performance versus fractional bandwidth for various beamformers and two setups including mmWave with $f_{0}=28$ GHz and $N=64$ antennas, and sub-THz with $f_{0}=140$ GHz and $N=256$ antennas. We consider $M=16$, $\tilde{\gamma}_{0}=0$ dB, and $\theta_{\mathrm{u}}=60^{\circ}$. Depending on the design parameters, the Sparse-TTD beamformer’s achieved performance lies at some point within the upper and lower bounds shown as the green-shaded region.
@@
(b): Cost trend versus operating fractional bandwidth. The regions where Non-TTD-WBBG, Sparse-TTD-NBBG, and Full-TTD-NBBG beamformers are more cost-effective are highlighted in orange, blue, and green, respectively.
@@
As discussed earlier, the fractional bandwidth plays a pivotal role in shaping both the performance and cost of wideband hybrid beamformers, making the selection of an optimal architecture a non-trivial design challenge, In this section, we present a cost-performance analysis using simplified system models to provide insights into this trade-off. Fig. 4-a compares the performance of several hybrid beamforming architectures: Non-TTD wideband beam-gain (Non-TTD-WBBG), Non-TTD narrowband beam-gain (Non-TTD-NBBG), Full-TTD-NBBG, and Sparse-TTD-NBBG, all evaluated as functions of the fractional bandwidth $B/f_{0}$, where $B$ is the total bandwidth, and $f_{0}$ represents the central carrier frequency. Unlike the Non-TTD-WBBG, which explicitly broadens the beam across the entire spectrum, the other architectures optimize performance without explicitly considering inter-carrier frequency interactions or beam widening. Therefore, to distinguish these approaches against WBBG, we have used the suffix NBBG representing a narrowband beam-gain approach for these structures. In the numerical results, we considered two setups: mmWave with $f_{0}=28$ GHz and $N=64$ antennas, and sub-THz with $f_{0}=140$ GHz and $N=256$ antennas. For the sub-THz, the attenuation by atmospheric gases was modeled according to ITU-R Recommendation P.676 with ambient temperature $T=15^{\circ}$C, dry-air pressure $P=101300$ Pa, and water-vapor density $W=7.5,\mathrm{g/m^{3}}$. To simplify the analysis, we ignore the bandwidth-dependent performance degradation of TTDs, which is commonly negligible in practice. While the Full-TTD-NBBG architecture can, in theory, eliminate beam squint entirely, making its performance nearly independent of the fractional bandwidth, the practical limitations of wideband PSs in Sparse-TTD and Non-TTD architectures introduce performance degradation as the operating bandwidth increases. To quantify this, we consider a simplified performance model. Consider a ULA transmitter having $N$ antennas with half-wavelength interelement spacing for the central carrier frequency $f_{0}$, transmitting over $2M+1$ uniformly spaced carrier frequencies across bandwidth $B$. For simplicity, a single antenna receiver is considered. Let the fractional baseband carrier vector be defined as $\boldsymbol{b}=\left[-\frac{B}{2f_{0}},\ldots,\frac{B}{2f_{0}}\right]^{\mathrm{T}}_{(2M+1)\times 1}$, and let $G_{\mathrm{NB}}(b,\theta_{\mathrm{u}})$ denote the array gain at fractional baseband carrier $b$ and observation angle $\theta_{\mathrm{u}}$.
First, we examine the Non-TTD narrowband beamforming performance gap. To do so, we start by formulating the performance of the ideal Full-TTD-NBBG represented as $R_{\mathrm{FN}}=\sum_{m=-M}^{M}\log\left({1+\tilde{\gamma}_{m}N\eta}\right)$ where $\tilde{\gamma}_{m}$ is the SNR at each antenna element at carrier mm [^1] Note that the Line-of-Sight (LoS) assumption, while simplifying the analysis across the three architectures, also represents the worst-case analysis in terms of beam squint mitigation. This is because the multipath propagation resulting from a non-LoS channel assumption provides a high angular diversity, which inherently decreases the amount of beam squint. , $N$ corresponds to the maximized array gain equal to the number of antenna elements, and $\eta<1$ captures the performance degradation due to hybrid beamforming relative to a fully digital implementation. In a similar way, if we replace the TTDs by PSs and implement the conventional narrowband beamforming, the performance of the Non-TTD-NBBG can be modeled as $R_{\mathrm{NN}}=\sum_{m=-M}^{M}\log\left(1+\tilde{\gamma}_{m}G(b_{m},\theta_{\mathrm{u}})\mathcal{E}(b_{m})\eta\right)$. Here, two factors reduce performance: (a) $\mathcal{E}(b)<1$, which is a decreasing function of $b$ that models the impact of frequency-dependent hardware impairments on the phase shifters evaluated at $f=f_{0}(1+b)$; and (b) $G(b_{m},\theta_{\mathrm{u}})\equiv G_{\mathrm{NB}}(b_{m},\theta_{\mathrm{u}})$ which represents the narrow-band beam gain for each carrier $b_{m}$ is lower than the optimal array gain $N$ due to the beam squint. This can be approximated by the array gain of a fully digital beamformer with ideal PSs, given by $G_{\mathrm{NB}}(b,\theta_{\mathrm{u}})\approx\left|\frac{\sin(N\pi\Delta_{b})}{\sin(\pi\Delta_{b})}\right|$, where $\Delta_{b}=\frac{b}{2}\sin(\theta_{\mathrm{u}})$. Finally, the performance gap between the Full-TTD-NBBG and Non-TTD-NBBG denoted by $\Delta R_{\mathrm{FN-NN}}=R_{\mathrm{FN}}-R_{\mathrm{NN}}$ is represented by
@@
$\Delta R_{\mathrm{FN-NN}}=\sum_{m=-M}^{M}\log\left(\frac{1+\tilde{\gamma}_{m}N\eta}{1+\tilde{\gamma}_{m}G(b_{m},\theta_{\mathrm{u}})\mathcal{E}(b_{m})\eta}\right).$ (2)
The normalized performance gap corresponding to $\Delta R_{\mathrm{FN-NN}}$ is shown in Fig. 4-a for the two simulation scenarios at mmWave and THz frequency bands, where a simple linear model is used for the $\mathcal{E}(b)$, with unity gain at $b=0$ and a maximum degradation of 6 dB at $b=\pm\frac{B}{2f_{0}}$. It is observed that the gap becomes significantly large for $B/f_{0}>0.05$, rendering this architecture suitable only for narrowband scenarios.
@@
The Non-TTD-WBBG beamformer mitigates this gap by optimizing a wideband beam pattern $G_{\mathrm{WB}}(b_{m},\theta_{\mathrm{u}})$ that remains approximately flat across all carriers $b_{m}$ via a max-min beam gain optimization. This optimization is a complex non-convex problem which can be approximately solved using convexification techniques, such as those in [1]. Once the beam pattern is optimized through calculating the optimal value of $G_{\mathrm{WB}}(b_{m},\theta_{\mathrm{u}}),\forall m$, the corresponding performance gap between the Full-TTD-NBBG and Non-TTD-WBBG is obtained as $\Delta R_{\mathrm{FN-NW}}=R_{\mathrm{FN}}-R_{\mathrm{NW}}$ obtained from (2) where $G(b_{m},\theta_{\mathrm{u}})=G_{\mathrm{WB}}(b_{m},\theta_{\mathrm{u}})$. As depicted in Fig. 4-a, the Non-TTD-WBBG architecture substantially reduces the performance gap compared to the Non-TTD-NBBG approach. However, the gap increases with fractional bandwidth due to two main factors: the degradation of wideband beam gain caused by the increased angular coverage required to accommodate squinted beams, and the worsening hardware impairments of PSs at higher fractional bandwidths. While this work does not explicitly model all hardware non-idealities to maintain a tractable and insightful comparison of the architectures, various impairments affect each structure to different extents in practice. For instance, phase noise predominantly impacts phase-shifter-based architectures (Non-TTD-NBBG and Sparse-TTD), especially at high fractional bandwidths. Other impairments, such as thermal drift and insertion loss, influence both PSs and TTD units, and therefore affect all three architectures to varying degrees.
@@
Now we discuss the performance of the Sparse-TTD-NBBG beamformer. Depending on the structure of the hybrid beamformer and the number of RF chains and TTDs employed for the Sparse-TTD architecture, the performance can be analyzed to lie between two bounds. The lower bound corresponds to the case where no TTDs are used, which can be approximated by the performance of the Non-TTD-WBBG architecture. The upper bound is achieved when the maximum number of TTDs (i.e., $N$) is utilized together with $N$ available PSs in a fully digital beamformer structure, under the assumption that PSs are still present, unlike in the Full-TTD configuration. This corresponds to allocating each antenna element a dedicated TTD cascaded with a PS. Since the cascade of a TTD and a PS is functionally equivalent to a TTD, the performance of this configuration is similar to that of Full-TTD, but with additional degradation due to the bandwidth-dependent impairments of the PSs. Accordingly, the performance upper bound can be approximated as $\sum_{m=-M}^{M}\log\left(1+\tilde{\gamma}_{m}N\mathcal{E}(b_{m})\right)$. Depending on the design parameters, the actual performance of the Sparse-TTD-NBBG beamformer will lie somewhere between these bounds, as illustrated by the green-shaded region in Fig. 4-a. Furthermore, the performance gap between the sub-THz and mmWave cases is significantly larger for the Non-TTD-NBBG scenario than that related to the Non-TTD-WBBG scenario. The reason is that at sub-THz frequencies, the larger number of array elements yields much sharper narrowband beams, which increases the likelihood of beam misalignment and thus degrades performance (except for the central carrier which experiences no squint). In contrast, in WBBG the beams are deliberately broadened, which mitigates misalignment and reduces the sensitivity of performance to array size differences between the mmWave and sub-THz regimes.
@@
Beyond performance, practical deployment also necessitates consideration of implementation cost. Due to the wide range of parameters influencing cost, the literature lacks a definitive comparison of the cost across different wideband hybrid beamforming architectures. However, estimated comparisons can be made based on qualitative cost considerations and implementation assumptions. Assuming similar baseband and RF chains across all architectures represented in Fig. 1, the primary cost differences arise from the use of TTD units and wideband PSs. First, we compare the costs of the Non-TTD-NBBG and Full-TTD-NBBG architectures, represented by the orange and green curves in Fig. 4-b. At small bandwidth, the Non-TTD-NBBG is cheaper; however, its cost grows sharply with bandwidth due to the complexity of wideband phase shifters (multi-section cascaded architecture of wideband PSs [2], complex structures for wideband NGD compensation [8], and several other practical challenges), eventually surpassing the Full-TTD whose TTDs cost are less sensitive to the bandwidth increase. We now elaborate on why the cost of the Sparse-TTD architecture (blue curve) falls between those of the other architectures for relatively low (below a certain threshold, $\mathrm{th}_{1}$) and very high (above another threshold, $\mathrm{th}_{2}$) fractional bandwidths, as observed in the figure. For narrowband operation ($B/f_{0}<\mathrm{th}_{1}$), the high cost of TTDs dominates. Consequently, the Full-TTD, Sparse-TTD, and Non-TTD correspond to the highest, intermediate, and lowest costs, respectively, reflecting their use of the largest, moderate, and zero number of TTDs. For ultra-wideband operation ($B/f_{0}>\mathrm{th}_{2}$), the escalating cost of wideband phase shifters dominates, reversing the order: Non-TTD becomes most expensive, Full-TTD least, and Sparse-TTD again lies in between. Among the studied architectures, Non-TTD-NBBG offers the lowest cost but suffers from severe performance degradation at higher fractional bandwidths, limiting its suitability to narrowband systems. The cost trend with respect to the desired fractional bandwidth for the other beamforming architectures is schematically illustrated in Fig. 4-b. Sparse-TTD architecture commonly incurs higher costs than the Non-TTD counterparts due to the additional expense of TTD units. However, this cost gap narrows as the fractional bandwidth increases, since wideband PSs, employed in both Sparse-TTD and Non-TTD-WBBG architectures, become increasingly expensive at higher bandwidths, with their cost growth rate eventually surpassing that of TTDs. More specifically, in ultra-wideband scenarios, the cost of Sparse-TTD and Non-TTD-WBBG architectures may ultimately exceed that of the Full-TTD as illustrated in Fig. 4-b. This is due to the increased complexity of compensating components (NGDs) and calibration mechanisms, required to maintain constant phase for wideband PSs at extreme bandwidths, as discussed in Section III.
@@
Finally, an overall cost-performance analysis can be conducted to guide the selection of the optimal architecture as illustrated in Fig. 5. Based on the preceding discussion, the Non-TTD-NBBG and Full-TTD-NBBG hybrid beamformers clearly emerge as the preferred choices for narrowband and ultra-wideband scenarios, respectively. More specifically, the Full-TTD-NBBG beamformer is the preferred option in ultra-wideband scenarios in terms of both performance and cost as shown in Figs. 4-a and 4-b, respectively, and highlighted in Fig. 5. In contrast, the preferred option for wideband scenarios can be Sparse-TTD-NBBG or Non-TTD-WBBG; however, the decision between the optimal architecture is more nuanced, as it depends on specific design parameters and the relative emphasis placed on performance versus implementation cost.